A laser is used to produce young fringes with slits separated by 0.50mm. The screen is 1.0m

     from the slits and 10 fringe separations occupy 12.5mm. What is the wavelength of the laser light?

     State the advantage of using light from a laser. 


1
Expert's answer
2021-03-14T19:41:25-0400

Given,

Slit separation "(d)=0.50mm"

"(d)=0.5\times 10^{-3}m"

Screen distance "(D)=1.0 m"

Number of fringe "(n)=10"

Fringe width "(w)=\frac{12.5mm}{10}=1.25mm=1.25\times 10^{-3}m"

"\lambda = \frac{wd}{D}"

"\Rightarrow \lambda = \frac{1.25\times 10^{-3}\times 0.5\times 10^{-3}m}{1.0 m}"

"\lambda= 625nm"

The diffraction patter in laser light is very less as compared to normal rays.